what is y in 6x-7y=10 if your tring to write it out in y=mx+b form
Answer by dpaulin(12)

6x-7y=10 subtract 6 from both sides of the equation
-7y=10-6x divide by -7 on both sides again to get y by itself
y = -10/7 + 6/7x put in proper form (be cautious with your signs)
y = 6/7x - 10/7